OpenStudy (anonymous):

Passing through (-2,1) and parallel to the line whose equation is x-3y=7. write an equation for the line in slope intercept form. Thats all I need because i keep messing up

OpenStudy (anonymous):

If two lines are parallel what is the relationship of the slopes?

OpenStudy (anonymous):

for the slopes i've got y-1=1/3(x+2)

OpenStudy (anonymous):

oh i've got it now and slope intersept isy=1/3x+5/3

OpenStudy (anonymous):

Great! you got it.

OpenStudy (anonymous):

thanks